zarkam21:

Help please on this SAT question I am practicing, Do not understand it and am freaking out

pooja195:

Post the question please. :)

NelsonTA:

^

NelsonTA:

Ill try to help if I can :3

ThisGirlPretty:

@Hero get ready to work your magic

zarkam21:

1 attachment
zarkam21:

@Vocaloid @Hero

NelsonTA:

ohhhh well I should just leave I know nothing about math xD

pooja195:

You need to find the solution to the system of equations this means that you need to find the numerical value of a and the numerical value of b and then subtract b-a

zarkam21:

Yes I understand that

zarkam21:

(a,b) is equivalent to (x,y)

pooja195:

Yes, are you familiar with solving system of equations?

zarkam21:

Yes but the (b+a) and (b-a) is confusing me in the problem

pooja195:

Don't be confused by it, solve it as you normally would, distribute the 3 to (a+b)

pooja195:

\[1.5a−4.5b=3a+3b\] Subtract 3a from both sides: \[1.5a−4.5b+−3a=ou get3a+3b+−3a\] You get: \[−1.5a−4.5b=3b\] Add 4b to both sides: \[−1.5a−4.5b+4.5b=3b+4.5b\] You get: \[−1.5a=7.5b\] Divide both sides by -1.5: \[\frac{ -1.5a }{ -1.5}=\frac{ 7.5b }{ -1.5a }\]

zarkam21:

Ugh I dont know im ending up with a mess everytime i try it

mhchen:

Nvm, I see what pooja is leading to. Solve for a in terms of b in the first equation, then solve for b in terms of a in the first equation. Then you'll find out the only value that a and b could be.

pooja195:

a=-5b Sub -5b in place of a in: \[5.5b=−2.5a+5b\] \[5.5b=−2.5−5b+5b\] \[5.5b=17.5b\] Subtract -17.5b to both sides: \[5.5b−17.5b=17.5b-17.5b\] To get: \[-12b=0\] Divide both sides boy -12 to get: \[\huge~b=0\]

pooja195:

Does this make sense?

zarkam21:

I always struggle with linear equations and my SAT is coming up and I'm scared, but yeah does make sense

pooja195:

I took the SAT for times this year before getting a decent score, it's hard at first, but once you keep practicing you'll get it :) Khan academy is the way to go for practice! You're on the right path. To complete the problem, we had said that a=-5b We had gotten b=0 from the work previously done so : \[\huge~a=-5(0)=0\] So a=0 and b=0, what would our final answer be?

pooja195:

*4 times

Vocaloid:

pooja is doing well so far, I am just trying to think of a way to do this without as much arithmetic or distribution

pooja195:

^I was hoping there would be a shortcut of some sort because this was the SAT, you wouldn't have this much time to do this much work on the actual test

zarkam21:

If you guys figure out a shortcut please share :S

pooja195:

Also a tip for the SAT, in my 4 times taking the test I have realized that the numbers are usually small when it comes to the free response math portion. This does not mean that the answer can't be big but typically it is small. This problem is a perfect example because the answer would be 0 (sorry for the giveaway) . As you can see the answer is a small number

zarkam21:

Yes, that is true ... If there are any other tips anybody has about the SAT PLEASE PLEASE SHARE!

pooja195:

Also, consider looking at this https://questioncove.com/study#/updates/59f6639f7103626820a59c00 From @Vocaloid

pooja195:

You looked at it lol *face palm* I'll pass on more things that I can think of

Nnesha:

pretty sure this will be on the calculator portion of the test especially with these decimals.. after distribution you can use your calculator to find a and b values

zarkam21:

Yes @pooja195 I looked at that and @Vocaloid gave me many valuable tips which I am following now, I got the khan academy idea from her, otherwise I did not know how to start

Vocaloid:

if you re-arrange the equations so they're in the same format you get 1.5a - 4.5b = 3(a+b) -2.5a + 5.5b = 5(b-a)

pooja195:

@Nnesha On khan academy the question is shown as a free response

Vocaloid:

adding them together you get - a + b = b - a = 3(a+b) + 5(b-a)

zarkam21:

OKau and the 1.5 , -2.5 etc are ignored

Vocaloid:

you don't "ignore" them, it's just that the decimals cancel out and give you b - a on the left side

Vocaloid:

anyway, if I did my math correctly I get a = 7b and you could substitute this back into the original equation and solve for b I guess? :S

Vocaloid:

using this logic I also get b = 0 and a = 0

zarkam21:

@pooja195

zarkam21:

I got something completely different but I am very sure I am wrong

Vocaloid:

so I just picked one of the original equations and let a = 7b and solved for b

zarkam21:

And how did you get a=7b again

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid adding them together you get - a + b = b - a = 3(a+b) + 5(b-a) \(\color{#0cbb34}{\text{End of Quote}}\) comes from this

Shadow:

Don't know why that link was messing up c;

zarkam21:

OKay will make notes :S

Vocaloid:

I took that and then solved for a in terms of b

Vocaloid:

by expanding everything and putting a and b on separate sides of the equation

zarkam21:

ANd then the other numbers canceled out right

zarkam21:

we were basically left with the variables

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid if you re-arrange the equations so they're in the same format you get 1.5a - 4.5b = 3(a+b) -2.5a + 5.5b = 5(b-a) \(\color{#0cbb34}{\text{End of Quote}}\) I would recommend starting from here

zarkam21:

okay thanks

Vocaloid:

add the two equations together

zarkam21:

and that is what makes them cancel out then we are left with - a + b = b - a = 3(a+b) + 5(b-a)

Vocaloid:

yes

Vocaloid:

then you distribute everything and solve for a in terms of b

zarkam21:

right

zarkam21:

I need help with one more type of problem let me repost it if you don't mind